OpenStudy (anonymous):

(x-1)^2/9+(y+5)^2/16=1 what is major axis

OpenStudy (anonymous):

16x^2 + 25y^2 + 32x - 150y - 159 =0 {(4X)^2+2*(4X)*4+4^2}-4^2+{(5Y)^2-2*(5Y)*15+15^2}-15^2-159=0 (4X+4)^2 + (5Y-15)^2 = 400 16(X+1)^2 + 25(Y-3)^2 =400...DIVIDING BY 400 THROUGHOUT.. (X+1)^2/25 + (Y-3)^2/16 =1 (X+1)^2/5^2 + (Y-3)^2/4^2 =1 COMPARING WITH ABOVE STANDARD EQN. CENTRE IS (-1,3) ECCENTRICITY E IS {(25-16)/25}^0.5=3/5 FOCI ARE (-1+5*3/5 ,3) AND (-1-5*3/5,3)=(2,3) AND (-4,-3) MAJOR AXIS LENGTH = 2*5=10